OpenStudy (anonymous):

6. b)Given that p and q are roots of the quadratic equation x^2-6x+h=0,whereas 3p and 3q are roots of the quadratic equation 3x^2+wx-2=0 . Calculate the possible values of h and w .

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

\[x=3\pm\sqrt{9-h}\]is a start those are the zeros of the first one

OpenStudy (anonymous):

you can put say \(p=3+\sqrt{9-h}\) and \( q=3-\sqrt{9-h}\) and figure find a quadratic that has those solutions that might help

OpenStudy (anonymous):

i still don't get it. @satellite73